The half-life for a first-order reaction is 32s. What was the original concentration if, after 2.0 minutes, the reactant concentration is 0.062


1
Expert's answer
2021-06-14T13:38:09-0400

Half life = 32 s


use relation between rate constant and half life of 1st order reaction


k = (ln 2) / k


= 0.693/(half life)


= 0.693/(32)


= 2.166×10^-2 s-1


we have:


[A] = 0.062 M


t = 2.0 minutes = 2×60 = 120 s


k = 2.166×10^-2 s-1


use integrated rate law for 1st order reaction


ln[A] = ln[A]o - k×t


ln(0.062) = ln[A]o - 2.166×10^-2×120


ln[A]o = -2.780621 + 2.166×10^-2×120

ln[A]o = -0.181871

[A]o = 0.834 M

Answer: 0.834 M